# New Applications of Ergodic Theory to Sets of Differences

### v1

@@ 4.3 Extending Theorem 1 to Cancellative Amenable Semigroups
Lemma 4.17. If $G$ is a countably infinite (left) amenable cancellative semigroup and $\widetilde{G}$ is its group of quotients then $G$ is right thick in $\widetilde{G}$.
Proof.
Let $(F_{N})_{N=1}^{\infty}\subset G$ be a (left) Følner sequence in $G$, and hence a Følner sequence in $\widetilde{G}$ by Theorem 4.3. Let $F=\{g_{1},\dots,g_{k}\}$ be any finite subset of $\widetilde{G}$. Since $(F_{N})_{N=1}^{\infty}$ is a Følner sequence in $\widetilde{G}$,
$\lim_{N\rightarrow\infty}\frac{\left|F_{N}\cap g_{1}^{-1}F_{N}\cap g_{2}^{-1}F_{N}\cap\dots\cap g_{n}^{-1}F_{N}\right|}{|F_{N}|}=1$
This means there exists $r\in\mathbb{N}$ such that $F_{r}\cap g_{1}^{-1}F_{r}\cap g_{2}^{-1}F_{r}\cap\dots\cap g_{k}^{-1}F_{r}\neq\emptyset$. Take any element $g\in F_{r}\cap g_{1}^{-1}F_{r}\cap g_{2}^{-1}F_{r}\cap\dots\cap g_{k}^{-1}F_{r}$ then $Fg\subset F_{r}$. $F$ was an arbitrary finite set, so $\bigcup_{N=1}^{\infty}F_{N}$ is right thick in $\widetilde{G}$. Since $(F_{N})_{N=1}^{\infty}$ is a Følner sequence in $G$, it follows that $\bigcup_{N=1}^{\infty}F_{N}\subset G$, and hence $G$ is right thick in $\widetilde{G}$. ∎
Remark 4.18. If $S\subset G$ is right thick in $G$ then $S$ is right thick in $\widetilde{G}$. Indeed, for any finite subset $F$ of $\widetilde{G}$ there exists a $g_{1}\in\widetilde{G}$ such that $Fg_{1}\subset G$ by Lemma 4.3. Since $Fg_{1}$ is a finite subset of $G$ there exists $g_{2}\in G$ such that $F(g_{1}g_{2})\subset S$.
To obtain a variant of Theorem 4.2 for semigroups we need to slightly modify the definition of tempered Følner sequences, but first we will need to introduce some notation. Let $G$ be a semigroup. For $B\subseteq G$, define
$A^{-1}B=\bigcup_{a\in A}a^{-1}B$
where $a^{-1}B=\{h\in G:ah\in B\}$.
Definition 4.19. A sequence $(F_{n})_{n=1}^{\infty}$ of finite subsets of $G$ is (left) tempered if for some $C>0$
@@
Theorem 4.21. Let $G$ be a countably infinite amenable cancellative semigroup and $\widetilde{G}$ be its group of quotients. Let $(F_{n})_{n=1}^{\infty}$ be a Følner sequence in $G$. Then the sequence $(F_{n})_{n=1}^{\infty}$ is a tempered Følner sequence in $G$ iff it is a tempered Følner sequence in $\widetilde{G}$
Proof.
Let $\varphi\colon G\to\widetilde{G}$ be an embedding of $G$ into its group of quotients $\widetilde{G}$. We wish to show that the sequence $(F_{n})_{n=1}^{\infty}$ is a tempered Følner sequence in $G$ if and only if $(\varphi(F_{n}))_{n=1}^{\infty}$ is a tempered Følner sequence in $\widetilde{G}$. Note that for all $a\in G$ and for $n\in\mathbb{N}$ with $n>1$,
$\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a\subseteq\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})a.$
@@
$\left|\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a\right|\leq\left|\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})a\right|=\left|\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})\right|\leq C|\varphi(F_{n})|=C|F_{n}|.$
@@
Now note that, for $n\in\mathbb{N}$ with $n>1$, since $\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})$ is a finite subset of $\widetilde{G}$, by Corollary 4.3 there exists an $a_{n}\in G$ such that $\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})a_{n}\subseteq G$. This means
$\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})a_{n}=\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a_{n}.$
@@
$\left|\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a\right|=\left|\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})a\right|=\left|\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})\right|>C|\varphi(F_{n})|=C|F_{n}|$
This shows that $(F_{n})_{n=1}^{\infty}$ is not a tempered Følner sequence. ∎
@@
Lemma 4.22. Let $G$ be a countably infinite amenable cancellative semigroup, let $\mathbf{F}$ be a Følner sequence in $G$, and let $S\subseteq G$ with $\overline{d}_{\mathbf{F}}(S)>0$. Let $E=\Delta_{1}(S)$. Then there exist $b_{1},\ldots,b_{\ell}\in G$ with $\ell\leq 1/\overline{d}_{\mathbf{F}}(S)$ for which $Eb_{1}^{-1}\cup\cdots\cup Eb_{\ell}^{-1}=G$.
Proof.
First, by Theorem 4.3, $G$ is a subsemigroup of its group of quotients $\widetilde{G}$ and $\mathbf{F}$ is a Følner sequence in $\widetilde{G}$. For purpose of clarity we will write $E_{G}$ and $\widetilde{E_{G}}$ to denote $\Delta_{1}(S)\cap G$ and $\Delta_{1}(S)\cap\widetilde{G}$, respectively, so
$\quad E_{G}=\{g\in G:S\cap g^{-1}S\neq\emptyset\}\quad\text{and}\quad\widetilde{E_{G}}=\{g\in\widetilde{G}:S\cap g^{-1}S\neq\emptyset\}.$
By Theorem 4.2 there are $a_{1},\ldots,a_{k}\in\widetilde{G}$ with $k\leq 1/\overline{d}_{\mathbf{F}}(S)$ such that $\widetilde{E_{G}}a_{1}^{-1}\cup\cdots\cup\widetilde{E_{G}}a_{k}^{-1}=\widetilde{G}$. A look at the proof of Theorem 4.2 reveals that the set $\{a_{1},\ldots,a_{k}\}$ has an additional property: $a_{i}F\cap a_{j}F=\emptyset$ for all finite $F\subseteq S$ and $i\neq j$. In fact $\{a_{1},\ldots,a_{k}\}$ is maximal with respect to this property in the sense that for any set $A\subseteq\widetilde{G}$ with more than $k$ elements there are $a,b\in A$ such that $aF\cap bF\neq\emptyset$ for some finite $F\subseteq S$. Therefore one may choose a subset $\{b_{1},\ldots,b_{\ell}\}$ of $G$ where $\ell\leq k$ such that if $g\in\widetilde{G}$, $g\notin\{b_{1},\ldots,b_{\ell}\}$, then there is some finite set $F\subseteq S$ and $1\leq i\leq\ell$ such that $gF\cap b_{i}F\neq\emptyset$. Note that $\ell\leq k\leq 1/\overline{d}_{\mathbf{F}}(S)$.
We now show that $E_{G}b_{1}^{-1}\cup\cdots\cup E_{G}b_{\ell}^{-1}=G$, as desired. Let $g\in G$. Then $g$ has some inverse $g^{-1}$ in the group $\widetilde{G}$. By the property of the set $\{b_{1},\ldots,b_{\ell}\}$ there is a finite set $F\subseteq S$ such that $g^{-1}F\cap b_{i}F\neq\emptyset$ for some $1\leq i\leq\ell$, so there is some $v\in F$ such that $b_{i}v\in g^{-1}F$. Hence $gb_{i}v=u$ for some $u\in F$. We have $v\in S$ and $gb_{i}v=u\in S$, so $v\in S\cap(gb_{i})^{-1}S$, thus $gb_{i}\in E_{G}$, therefore $g\in E_{G}b_{i}^{-1}$. ∎
@@
Theorem 4.23. Let $G$ be a countably infinite amenable cancellative semigroup and let $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ be Følner sequences in $G$ such that $\mathbf{G}$ is tempered. If $S_{1},\ldots,S_{k}\subseteq G$ satisfy $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$, then there is $S\subseteq G$ such that $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D:=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in G$ where
@@
such that $\bigcup_{i=1}^{\ell}(Dm_{i}^{-1})=G$.
Proof.
Let $\widetilde{G}$ be the group of quotients of $G$. Since in this proof we will be working simultaneously with subsets of $G$ and $\widetilde{G}$, it will be convenient to slightly modify the notation introduced in Definition 4.2. Namely for any $A\subset G$, $B\subset\widetilde{G}$, and Følner sequence $\mathbf{F}$ in $G$ (considered, when convenient, as a Følner sequence in $\widetilde{G}$) we will write
$\Delta_{1}(A,G)=\{g\in G:A\cap g^{-1}A\neq\emptyset\}\quad\text{and}\quad\Delta_{2}(\mathbf{F},A,G)=\{g\in G:\overline{d}_{\mathbf{F}}(A\cap g^{-1}A)>0\}.$
$\Delta_{1}(B,\widetilde{G})=\{g\in\widetilde{G}:B\cap g^{-1}B\neq\emptyset\}\quad\text{and}\quad\widetilde{\Delta_{2}}(\mathbf{F},B,\widetilde{G})=\{g\in\widetilde{G}:\overline{d}_{\mathbf{F}}(B\cap g^{-1}B)>0\}.$
Theorem 4.3 implies that $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ are Følner sequences in $\widetilde{G}$. Additionally, Theorem 4.3 shows $\mathbf{G}$ is tempered in $\widetilde{G}$. So by Theorem 4.2, there exists $S^{\prime}\subset\widetilde{G}$ such that $d_{\mathbf{G}}(S^{\prime})\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S^{\prime},\widetilde{G})\subset\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i},\widetilde{G})$. Let $S\vcentcolon=S^{\prime}\cap G$. Note by Remark 4.3 that
$d_{\mathbf{G}}(S)=d_{\mathbf{G}}(S^{\prime})\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$
and
$\Delta_{1}(S,G)\subset\Delta_{1}(S^{\prime},G)\subset\Delta_{1}(S^{\prime},\widetilde{G})\subset\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i},\widetilde{G})$
This implies
$\Delta_{1}(S,G)\subset\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i},\widetilde{G})\cap G=D.$
@@
such that $(Dm_{1}^{-1})\cup\cdots\cup(Dm_{k}^{-1})=G$. ∎

### v2

@@ 4.2 Extending Theorem 1 to amenable groups
Lemma 4.18. If $G$ is a countably infinite (left) amenable cancellative semigroup and $\widetilde{G}$ is its group of quotients, then for any finite $F\subseteq\widetilde{G}$, there exists $g\in G$ such that $Fg\subseteq G$. In particular, $G$ is right thick in $\widetilde{G}$.
Proof. Let $(F_{N})_{N=1}^{\infty}$ be a (left) Følner sequence in $G$, and hence a Følner sequence in $\widetilde{G}$ by Theorem 4.3. Let $F=\{g_{1},\dots,g_{k}\}$ be any finite subset of $\widetilde{G}$. Since $(F_{N})_{N=1}^{\infty}$ is a Følner sequence in $\widetilde{G}$,
$\lim_{N\rightarrow\infty}\frac{\left|F_{N}\cap g_{1}^{-1}F_{N}\cap g_{2}^{-1}F_{N}\cap\dots\cap g_{n}^{-1}F_{N}\right|}{|F_{N}|}=1.$
This means there exists $r\in\mathbb{N}$ such that $F_{r}\cap g_{1}^{-1}F_{r}\cap g_{2}^{-1}F_{r}\cap\dots\cap g_{k}^{-1}F_{r}\neq\emptyset$. Now, for any element $g\in F_{r}\cap g_{1}^{-1}F_{r}\cap g_{2}^{-1}F_{r}\cap\dots\cap g_{k}^{-1}F_{r}$, we have $Fg\subseteq F_{r}\subseteq G$. Since $F$ was an arbitrary finite subset of $\widetilde{G}$, it follows that $\bigcup_{N=1}^{\infty}F_{N}$ is right thick in $\widetilde{G}$. Since $(F_{N})_{N=1}^{\infty}$ is a Følner sequence in $G$, it follows that $\bigcup_{N=1}^{\infty}F_{N}\subseteq G$, and hence $G$ is right thick in $\widetilde{G}$. $\square$
Remark 4.19. If $S\subseteq G$ is right thick in $G$, then $S$ is right thick in $\widetilde{G}$. Indeed, for any finite subset $F$ of $\widetilde{G}$ there exists a $g_{1}\in\widetilde{G}$ such that $Fg_{1}\subseteq G$ by Lemma 4.3. Since $Fg_{1}$ is a finite subset of $G$ there exists $g_{2}\in G$ such that $Fg_{1}g_{2}\subseteq S$.
To obtain a variant of Theorem 4.2 for semigroups, one needs to slightly modify the definition of tempered Følner sequences, but first we will need to introduce some notation. Let $G$ be a semigroup. For $B\subseteq G$, define
$A^{-1}B=\bigcup_{a\in A}a^{-1}B,$ (4.15)
where $a^{-1}B=\{h\in G:ah\in B\}$. Note that when $G$ is a group, $A^{-1}B=\{a^{-1}b:a\in A,b\in B\}$.
Definition 4.20. A sequence $(F_{n})_{n=1}^{\infty}$ of finite subsets of a semigroup $G$ is (left) tempered if for some $C>0$
@@
Theorem 4.22. Let $G$ be a countably infinite amenable cancellative semigroup and $\widetilde{G}$ be its group of quotients. Let $(F_{n})_{n=1}^{\infty}$ be a Følner sequence in $G$. Then the sequence $(F_{n})_{n=1}^{\infty}$ is a tempered Følner sequence in $G$ iff it is a tempered Følner sequence in $\widetilde{G}$.
Proof. The fact that $(F_{n})_{n=1}^{\infty}$ is a Følner sequence in $G$ iff it is a Følner sequence in $\widetilde{G}$ follows from Theorem 4.3. Let $\varphi\colon G\to\widetilde{G}$ be an embedding of $G$ into its group of quotients $\widetilde{G}$. We wish to show that the sequence $(F_{n})_{n=1}^{\infty}$ is a tempered Følner sequence in $G$ if and only if $(\varphi(F_{n}))_{n=1}^{\infty}$ is a tempered Følner sequence in $\widetilde{G}$. Note that for all $a\in G$ and for $n\in\mathbb{N}$ with $n>1$,
$\varphi\left(\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a\right)\subseteq\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})\varphi(a).$
@@
$\left|\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a\right|=\left|\varphi\left(\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a\right)\right|$ $\leq\left|\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})\varphi(a)\right|$
$=\left|\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})\right|\leq C|\varphi(F_{n})|=C|F_{n}|.$
@@
Now note that, for $n\in\mathbb{N}$ with $n>1$, since $\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})$ is a finite subset of $\widetilde{G}$, by Lemma 4.3 there exists an $a_{n}\in G$ such that $\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})\varphi(a_{n})\subseteq\varphi(G)$. This means
$\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})\varphi(a_{n})=\varphi\left(\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a_{n}\right).$
@@
$\left|\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a_{n}\right|=\left|\varphi\left(\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}a_{n}\right)\right|$ $=\left|\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})\varphi(a_{n})\right|$
$=\left|\bigcup_{k=1}^{n-1}\varphi(F_{k})^{-1}\varphi(F_{n})\right|>C|\varphi(F_{n})|=C|F_{n}|$
This shows that $(F_{n})_{n=1}^{\infty}$ is not a tempered Følner sequence in $G$. $\square$
@@
Lemma 4.23. Let $G$ be a countably infinite amenable cancellative semigroup, let $\mathbf{F}$ be a Følner sequence in $G$, let $S\subseteq G$ with $\overline{d}_{\mathbf{F}}(S)>0$, and let $E=\Delta_{1}(S)$. Then there exist $b_{1},\ldots,b_{\ell}\in G$ with $\ell\leq 1/\overline{d}_{\mathbf{F}}(S)$ for which $Eb_{1}^{-1}\cup\cdots\cup Eb_{\ell}^{-1}=G$.
Proof. Let $\widetilde{E}=\{g\in\widetilde{G}:S\cap g^{-1}S\neq\emptyset\}$. Now, since F is a Følner sequence in $\widetilde{G}$ by Theorem 4.3, it follows from Theorem 4.2 that there exists $k\leq 1/\overline{d}_{\textbf{F}}(S)$ (here, we regard $S$ as a subset of $\widetilde{G}$) and $a_{1},...,a_{k}\in\widetilde{G}$ such that $\widetilde{E}a_{1}^{-1}\cup\cdots\cup\widetilde{E}a_{k}^{-1}=\widetilde{G}$. A look at the proof of Theorem 4.2 reveals that the set $\{a_{1},\ldots,a_{k}\}$ has an additional property: $a_{i}F\cap a_{j}F=\emptyset$ for all finite $F\subseteq S$ and all $i\neq j$. In fact $\{a_{1},\ldots,a_{k}\}$ is maximal with respect to this property in the sense that for any set $A\subseteq\widetilde{G}$ with more than $k$ elements there are $a,b\in A$ such that $aF\cap bF\neq\emptyset$ for some finite $F\subseteq S$. Therefore one may choose a subset $\{b_{1},\ldots,b_{\ell}\}$ of $G$, where $\ell\leq k$, such that if $g\in\widetilde{G}\setminus\left\{b_{1},...,b_{\ell}\right\}$, then there is some finite set $F\subseteq S$ and $1\leq i\leq\ell$ such that $gF\cap b_{i}F\neq\emptyset$. Note that $\ell\leq k\leq 1/\overline{d}_{\mathbf{F}}(S)$.
We now show that $Eb_{1}^{-1}\cup\cdots\cup Eb_{\ell}^{-1}=G$, as desired. Let $g\in G$. Then $g$ has some inverse $g^{-1}$ in the group $\widetilde{G}$. By the property of the set $\{b_{1},\ldots,b_{\ell}\}$ there is a finite set $F\subseteq S$ such that $g^{-1}F\cap b_{i}F\neq\emptyset$ for some $1\leq i\leq\ell$, so there exist $u,v\in F$ such that $gb_{i}v=u$. We have $v\in S$ and $gb_{i}v=u\in S$, so $v\in S\cap(gb_{i})^{-1}S$, thus $gb_{i}\in E$, meaning $g\in Eb_{i}^{-1}$. $\square$
@@
Theorem 4.24. Let $G$ be a countably infinite amenable cancellative semigroup and let $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ be Følner sequences in $G$ such that $\mathbf{G}$ is tempered. If $S_{1},\ldots,S_{k}\subseteq G$ satisfy $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$, then there is $S\subseteq G$ such that $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D:=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in G$, where
@@
such that $\bigcup_{i=1}^{\ell}(Dm_{i}^{-1})=G$. If in addition $G$ is a group, then $\bigcup_{i=1}^{\ell}(m_{i}D)=G$ as well.
Proof. Let $\widetilde{G}$ be the group of quotients of $G$. Since in this proof we will be working simultaneously with subsets of $G$ and $\widetilde{G}$, it will be convenient to slightly modify the notation introduced in Definition 4.2. Namely for any $A\subseteq G$, $B\subseteq\widetilde{G}$, and Følner sequence $\mathbf{F}$ in $G$ (considered, when convenient, as a Følner sequence in $\widetilde{G}$) we will write
$\Delta_{1}(A,G)=\{g\in G:A\cap g^{-1}A\neq\emptyset\}\quad\text{and}\quad\Delta_{2}(\mathbf{F},A,G)=\{g\in G:\overline{d}_{\mathbf{F}}(A\cap g^{-1}A)>0\},$
$\Delta_{1}(B,\widetilde{G})=\{g\in\widetilde{G}:B\cap g^{-1}B\neq\emptyset\}\quad\text{and}\quad\Delta_{2}(\mathbf{F},B,\widetilde{G})=\{g\in\widetilde{G}:\overline{d}_{\mathbf{F}}(B\cap g^{-1}B)>0\}.$
Theorem 4.3 implies that $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ are Følner sequences in $\widetilde{G}$. Additionally, Theorem 4.3 shows $\mathbf{G}$ is tempered in $\widetilde{G}$. So by Theorem 4.2, there exists $S^{\prime}\subseteq\widetilde{G}$ such that $d_{\mathbf{G}}(S^{\prime})\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S^{\prime},\widetilde{G})\subseteq\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i},\widetilde{G})$. Let $S\vcentcolon=S^{\prime}\cap G$. Note that by Remark 4.3
$d_{\mathbf{G}}(S)=d_{\mathbf{G}}(S^{\prime})\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i}).$
Also observe that,
$\Delta_{1}(S,G)\subseteq\Delta_{1}(S^{\prime},G)\subseteq\Delta_{1}(S^{\prime},\widetilde{G})\subseteq\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i},\widetilde{G})$
which implies
$\Delta_{1}(S,G)\subseteq\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i},\widetilde{G})\cap G=D.$
@@
such that $(Dm_{1}^{-1})\cup\cdots\cup(Dm_{k}^{-1})=G$. If $G$ is a group, then the fact that $\bigcup_{i=1}^{\ell}(m_{i}D)=G$ follows from Theorem 4.2. $\square$